OpenStudy (anonymous):

Find the solution to the equations. 3x - y = -4 x + y = 0

OpenStudy (anonymous):

x=-y -3y -y = -4 -4y = -4 y=1 x=-1

OpenStudy (anonymous):

y=3x y=4x x=0 y=0